OpenStudy (anonymous):

Find the derivative of f(x) = negative 6 divided by x at x = 12 and Find the derivative of f(x) = 6x + 2 at x = 1.

OpenStudy (freckles):

\[f(x)=\frac{-6}{x}\] are you suppose to find f' using the definition of derivative?

OpenStudy (anonymous):

Do I plug it in

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

Im just confused since my textbook teaches it like that but doesnt really explain

OpenStudy (freckles):

\[f(x)=\frac{-6}{x} \\ f(x+h)=\frac{-6}{x+h}\]

OpenStudy (freckles):

plug those into the definition above

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{\frac{-6}{x+h}-\frac{-6}{x}}{h} =\lim_{h \rightarrow 0}\frac{1}{h}(\frac{-6}{x+h}-\frac{-6}{x}) \] combine those fractions inside the ( )

OpenStudy (freckles):

we are trying to get a factor of h in the numerator that will cancel with that factor of h in the denominator

OpenStudy (anonymous):

how do i know to put it under x+h

OpenStudy (freckles):

what?

OpenStudy (anonymous):

Nevermind

OpenStudy (freckles):

I'm using the definition of derivative \[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] all I did was replace f(x+h) and f(x)

OpenStudy (freckles):

f(x)=-6/x so f(x+h)=-6/(x+h)

OpenStudy (freckles):

so I just put those two things in

OpenStudy (anonymous):

Oh ok I see

OpenStudy (freckles):

so you need to combine the fractions in the ( )

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{\frac{-6}{x+h}-\frac{-6}{x}}{h} =\lim_{h \rightarrow 0}\frac{1}{h}(\frac{-6}{x+h}-\frac{-6}{x}) \]

OpenStudy (freckles):

we are trying to get a factor of h on top so we can cancel the h on bottom

OpenStudy (anonymous):

So the h and x cancel

OpenStudy (anonymous):

not the x oops

OpenStudy (anonymous):

so 6/12?

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

no right

OpenStudy (freckles):

have you combined the fractions in the ( ) yet

OpenStudy (anonymous):

(-6/x)- (-6/x)?

OpenStudy (freckles):

what happen to the h?

OpenStudy (anonymous):

i thought i ws supposed to cancel it

OpenStudy (freckles):

you have -6/(x+h)-(-6/x)

OpenStudy (freckles):

with what?

OpenStudy (freckles):

h don't just go away

OpenStudy (freckles):

the whole reason we are combine the fractions is so we can cancel the factor h on the outside of the ( )

OpenStudy (anonymous):

so what is the valye of h

OpenStudy (anonymous):

is the answer 2

OpenStudy (freckles):

why won;t you try to combine the fractions?

OpenStudy (freckles):

h is a variable

OpenStudy (anonymous):

oh okay let me try

OpenStudy (anonymous):

umm -6/h plus 6

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{\frac{-6}{x+h}-\frac{-6}{x}}{h} =\lim_{h \rightarrow 0}\frac{1}{h}(\frac{-6}{x+h}-\frac{-6}{x}) \\ =\lim_{h \rightarrow 0} \frac{1}{h}(\frac{-6(x)-(-6)(x+h)}{(x)(x+h)}) \\ \]

OpenStudy (freckles):